OpenStudy (anonymous):

Find all solutions of the equation sec4x−2=0 The answer is A+k2 and B+k2 where k is any integer and 0

OpenStudy (anonymous):

\[\sec (4x)-2=0\] \[=1/\cos(4x)=2\] \[=>\cos(4x)=1/2\] \[4x=\cos^{-1}(1/2)\] \[4x=\cos^{-1}(0.5)+2k\pi,\ \ k\in \mathbb{Z}\] \[x=arcCos(0.5)/4 + k·\frac{ \pi }{ 2}, \ \ k\in \mathbb{Z}\] or X=15º (degree) +k 90º

OpenStudy (anonymous):

the reason to have the term 2kpi is because there are many solutions to Cos (x)=0.5 everyone is delayed 360º but i made a mistake, because Cos( X)= cos(-x)=0.5 And de delay is not anymore 360º now it is 180º or "pi" \[4x=\cos^{-1} (0.5) + k\pi\] \[X=\cos^{-1}(0.5)/4+k·\pi /4\]

OpenStudy (anonymous):

So A=arccos(0.5)/4 and B=PI/4??

OpenStudy (anonymous):

they ask you 2 solutions X1=A+K1, X2 =B+K2 then A = B = arccos(0.5)/4 But K1 = when K=1 and K2 = when K =2

OpenStudy (anonymous):

if i'm right, the solutions are: X1=arccos(0.5)/4 + pi/4 X2=arccos(0.5)/4 + pi/2 i don't know what 0<A<B<pi/2 means. i assume typing error maybe A and B are the solutions So A=arccos(0.5)/4+pi/4 B=the same.... you can choose any positive or negative value (integer) of k

OpenStudy (anonymous):

Okay, thank you soo much for all your help!